OpenStudy (anonymous):

What is the solution of the linear-quadratic system of equations? {y = x^2 + 5x - 3 {y - x = 2

OpenStudy (anonymous):

Use substitution. We know that: \(y = x+2\) Plug it in: \(x+2 = x^2+5x-3\) Do you know what to do from there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[x^2+4x-5 = 0\] \[(x-1)(x+5) = 0\] x = -5 or 1 Now you know \(x\) so solve for \(y\).

OpenStudy (anonymous):

Oh ok I see, one second

OpenStudy (anonymous):

So: y = -2 or 3 right?

OpenStudy (anonymous):

Yes, you are correct :-).